(5.2.28) Based on a​ poll, among adults who regret getting​ tattoos, ​19% say that they were too young when they got their tattoos. Assume that 5 adults who regret getting tattoos are randomly​ selected, and find the indicated probability.
A.Find the probability that none of the selected adults say that they were too young to get tattoos
1
Expert's answer
2020-12-03T11:16:42-0500

p = 19% = 0.19

n = 5

x = 0

Binomial probability:

"P(X=x)=C(n,x)\cdot p^x\cdot (1-p)^{n-x}"

"P(X=0)=\frac{5!}{0!5!}\cdot0.19^0\cdot0.81^5=0.35"


Answer: the probability that none of the selected adults say that they were too young to get tattoos is 0.35.